OpenStudy (anonymous):

PLEASE HELP!!!!!!! Let f(x) = 5x2 + 5x − 4. Find and simplify the following. f (4/a)

OpenStudy (anonymous):

\[f(\frac{ 4 }{ a }) = 5(\frac{ 4 }{ a })^{2}+5(\frac{ 4 }{ a })-4\]

OpenStudy (anonymous):

Can you solve it from that?

OpenStudy (anonymous):

no thats where im stuck

OpenStudy (anonymous):

Okay, be sure to square both the 4 and the a in the first term

OpenStudy (anonymous):

so you would end up with 5(16/a^2) +20/a - 4

OpenStudy (anonymous):

okay then what?

OpenStudy (anonymous):

Then you need to combine into a fraction by multiplying the fractions by the LCD so you'd have 5 x (16+20a)/(a^2)

OpenStudy (anonymous):

then you still have the -4 to deal with; so in the end you'll have 5 x (16+20a-4a^2)/a^2

OpenStudy (anonymous):

then simplify from there, let me know what you get

OpenStudy (anonymous):

80+100a-20a^2 / a^2

OpenStudy (anonymous):

are there any terms you can cancel?

OpenStudy (anonymous):

a^2 ?

OpenStudy (anonymous):

so would it be 80+100a-20 ?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes, one sec

OpenStudy (anonymous):

ok so the final answer is 100a+60?

OpenStudy (anonymous):

I'd leave it as the fraction: (-4a^2 + 20a +80)/a^2

OpenStudy (anonymous):

ok thanks